OpenStudy (anonymous):

During the first part of a trip, Tara's Honda traveled 120 miles at a certain speed. Tara then drove another 100 miles at a speed that was 10 mph slower. If the total time of Tara's trip was 4 hrs, what was her speed on each part of the trip?

OpenStudy (anonymous):

set up equation

OpenStudy (anonymous):

i need help setting up the equation. after that I can solve

OpenStudy (anonymous):

(120/v)+(100/(v-10)=4 (120v-1200+100v)/(v*v-10v)=4 220v-1200=4v*v-40v 55v-300=v*v-10v v*v-65v+300=0 (v-60)*(v-5)=0 then it will be 60 and it can not be 5